OpenStudy (anonymous):

Two shirts and one tie cost $42, but one shirt and two ties cost $39. What does one shirt cost?

OpenStudy (anonymous):

let x be the price of the shirt let y be the price of the tie 2 shirt and 1 tie cost $42 2x+1y=42 does this make sense?

OpenStudy (lgbasallote):

No.

OpenStudy (anonymous):

well i tried it and i got some crazy answer that was not an option it didnt work...

OpenStudy (anonymous):

Okay, if 2s + t = 42 and s + 2t = 39, then: s + 2t + 3 = 2s + t With me so far?

OpenStudy (anonymous):

From this, -s + t = -3, so t = s - 3

OpenStudy (anonymous):

:D Yes

OpenStudy (anonymous):

Substitute s - 3 in for t in one of your original formulas, and solve for s.

OpenStudy (anonymous):

wait in the equation s+2t+3=2s+t where did u get the three from?

OpenStudy (anonymous):

Sorry... You knew the following: 2s + t = 42 and s + 2t = 39 For the second formula, to make it equal 42, I added 3 on both sides, hence: s + 2t + 3 = 42 Then, I could equate the left side of formula 1 to the left side of formula 2, because 2s + t = 42 = s + 2t + 3

OpenStudy (anonymous):

I shouldn't say "you knew the following." It was intended as "the following was given." The first way sounds rude.

OpenStudy (anonymous):

I would just use the substitution method. Less thinking involved = less mistakes

OpenStudy (anonymous):

I would choose elimination any day...

OpenStudy (anonymous):

personal preference I guess

OpenStudy (anonymous):

@ChmE I've no idea what that is, but if you want to write out the steps I'm sure it would help either Jocelynna or someone else who comes along and sees this question.

OpenStudy (anonymous):

i tried substitution i got some like really high insane number susbstitution doesnt work in this case

OpenStudy (anonymous):

2x + y = 42 x + 2y = 39 2(39-2y) + y =42 78 - 4y + y = 42 -3y = -36 y = 12 2x+ 12 = 42 2x = 30 x = 15

OpenStudy (anonymous):

x are shirts y are ties

OpenStudy (anonymous):

@ChmE thanks so much your method was what im learning i just got minor algebra mistakes

OpenStudy (anonymous):

You're welcome. Just take your time at each step and you'll be fine. Good luck

OpenStudy (anonymous):

@geoffb thanks so much for the help, @Dido525 thanks!

OpenStudy (anonymous):

:)

OpenStudy (konradzuse):

iggy is correct.